# Double Yangians and quantum vertex algebras, I

### v1

@@ Abstract.
For any simply-laced generalized Cartan matrix $A$, we introduce an algebra $\widehat{\mathcal{DY}}(A)$, which is essentially the centrally extended double Yangian when $A$ is of finite type, and we give a new field (current) presentation of $\widehat{\mathcal{DY}}(A)$. Among the main results, for any $\ell\in{\mathbb{C}}$ we construct a universal vacuum $\widehat{\mathcal{DY}}(A)$-module $\mathcal{V}_{A}(\ell)$ of level $\ell$, prove that there exists a natural $\hbar$-adic weak quantum vertex algebra structure on $\mathcal{V}_{A}(\ell)$, and give an isomorphism between the category of restricted $\widehat{\mathcal{DY}}(A)$-modules of level $\ell$ and the category of $\mathcal{V}_{A}(\ell)$-modules.
@@ 1. Introduction
In a previous paper [JKLT2] joint with Jing and Tan, by using the smash product construction of nonlocal vertex algebras (see [Li4]), we constructed a family of $\hbar$-adic quantum vertex algebras by explicitly deforming lattice vertex algebras. In [Kon], mainly by deforming affine vertex algebras the first named author of this paper constructed a family of $\hbar$-adic quantum affine vertex algebras and their $\phi$-coordinated quasi modules.
In this current paper, we continue studying centrally extended double Yangians in the context of $\hbar$-adic quantum vertex algebras and their modules, using the generator-and-defining-relation approach. We first introduce an algebra $\widehat{\mathcal{DY}}(A)$ for any simply-laced generalized Cartan matrix $A$, which essentially coincides with the centrally extended double Yangian when $A$ is of finite type, and we then give a new field (current) presentation of $\widehat{\mathcal{DY}}(A)$. Among the main results, for any $\ell\in{\mathbb{C}}$, we construct a universal vacuum $\widehat{\mathcal{DY}}(A)$-module $\mathcal{V}_{A}(\ell)$ of level $\ell$, prove that there exists a natural $\hbar$-adic weak quantum vertex algebra structure on $\mathcal{V}_{A}(\ell)$, and give an isomorphism between the category of restricted $\widehat{\mathcal{DY}}(A)$-modules of level $\ell$ and the category of $\mathcal{V}_{A}(\ell)$-modules.
@@ 4. Vertex-operator Serre relation and iterate formula
Let $a(x),b(x)\in{\mathcal{E}}_{\hbar}(W)$. The relation
(4.7) $\sum_{\sigma\in S_{2}}\left(a(z_{\sigma(1)})a(z_{\sigma(2)})b(w)-2a(z_{\sigma(1)})b(w)a(z_{\sigma(2)})+b(w)a(z_{\sigma(1)})a(z_{\sigma(2)})\right)=0$
is referred to as the order-$2$ Serre relation for $(a(z_{1}),a(z_{2}),b(w))$.
The following is a key result (cf. [DK, Prop. 1]):
Proposition 4.5.
Let $a(x),b(x)\in{\mathcal{E}}_{\hbar}(W)$ such that
(4.8) $(z-w-2\hbar)a(z)a(w)$ $=(z-w+2\hbar)a(w)a(z),$
(4.9) $(z-w+\hbar)a(z)b(w)$ $=(z-w-\hbar)b(w)a(z).$
Then the order-$2$ Serre relation for $(a(z_{1}),a(z_{2}),b(w))$ holds if and only if
(4.10) $a(z)a(w)_{0}b(w)=a(w)_{0}b(w)a(z)\!\left(\frac{w-z-3\hbar}{w-z-\hbar}\right)\!,$
where
$a(w)_{0}b(w)=\operatorname{Res}_{x_{1}}\!\left(\!a(x_{1})b(w)-\left(\frac{w-x_{1}+\hbar}{w-x_{1}-\hbar}\right)b(w)a(x_{1})\!\right)\!.$
Proof.
We express all the terms in terms of $b(w)a(z_{1})a(z_{2})$. With commutation relation (4.9), by Lemma 4.3 we have
$a(x_{1})b(x)=\left(\frac{x-x_{1}+\hbar}{x-x_{1}-\hbar}\right)b(x)a(x_{1})+x_{1}^{-1}\delta\!\left(\frac{x-\hbar}{x_{1}}\right)\!a(x)_{0}b(x).$
Using this we get
$a(z_{1})a(z_{2})b(w)=\$$a(z_{1})b(w)a(z_{2})\!\left(\frac{w-z_{2}+\hbar}{w-z_{2}-\hbar}\right)\!+a(z_{1})a(w)_{0}b(w)z_{2}^{-1}\delta\!\left(\frac{w-\hbar}{z_{2}}\right)$
[…]
$+a(w)_{0}b(w)a(z_{2})z_{1}^{-1}\delta\!\left(\frac{w-\hbar}{z_{1}}\right)\!\left(\frac{w-z_{2}+\hbar}{w-z_{2}-\hbar}\right)$
$+a(z_{1})a(w)_{0}b(w)z_{2}^{-1}\delta\!\left(\frac{w-\hbar}{z_{2}}\right)$
and
$a(z_{1})b(w)a(z_{2})=b(w)a(z_{1})a(z_{2})\!\left(\frac{w-z_{1}+\hbar}{w-z_{1}-\hbar}\right)\!+a(w)_{0}b(w)a(z_{2})z_{1}^{-1}\delta\!\left(\frac{w-\hbar}{z_{1}}\right)\!.$
Note that
$\frac{w-z_{1}+\hbar}{w-z_{1}-\hbar}\cdot\frac{w-z_{2}+\hbar}{w-z_{2}-\hbar}-2\left(\frac{w-z_{1}+\hbar}{w-z_{1}-\hbar}\right)+1=\frac{2\hbar(z_{2}-z_{1}+2\hbar)}{(w-z_{1}-\hbar)(w-z_{2}-\hbar)},$
$\frac{w-z_{2}+\hbar}{w-z_{2}-\hbar}-2=\frac{z_{2}-w+3\hbar}{w-z_{2}-\hbar}.$
Then we get
(4.11) $a(z_{1})a(z_{2})b(w)-2a(z_{1})b(w)a(z_{2})+b(w)a(z_{1})a(z_{2})$
$=\$$\frac{2\hbar(z_{2}-z_{1}+2\hbar)}{(w-z_{1}-\hbar)(w-z_{2}-\hbar)}b(w)a(z_{1})a(z_{2})$
$+\left(\frac{z_{2}-w+3\hbar}{w-z_{2}-\hbar}\right)a(w)_{0}b(w)a(z_{2})z_{1}^{-1}\delta\!\left(\frac{w-\hbar}{z_{1}}\right)$
$+a(z_{1})a(w)_{0}b(w)z_{2}^{-1}\delta\!\left(\frac{w-\hbar}{z_{2}}\right)\!.$
Noticing that
$2\hbar(z_{2}-z_{1}+2\hbar)b(w)a(z_{1})a(z_{2})+2\hbar(z_{1}-z_{2}+2\hbar)b(w)a(z_{2})a(z_{1})=0,$
we see that the order-$2$ Serre relation for $(a(z_{1}),a(z_{2}),b(w))$ holds if and only if
(4.12) $\left(\left(\frac{z_{2}-w+3\hbar}{w-z_{2}-\hbar}\right)a(w)_{0}b(w)a(z_{2})+a(z_{2})a(w)_{0}b(w)\!\right)\!z_{1}^{-1}\delta\!\left(\frac{w-\hbar}{z_{1}}\right)$
$=\$$-\!\left(\left(\frac{z_{1}-w+3\hbar}{w-z_{1}-\hbar}\right)a(w)_{0}b(w)a(z_{1})+a(z_{1})a(w)_{0}b(w)\!\right)\!z_{2}^{-1}\delta\!\left(\frac{w-\hbar}{z_{2}}\right)\!.$
Notice that multiplying (4.12) by $(w-z_{1}-\hbar)$ and then applying $\operatorname{Res}_{z_{2}}$ gives
(4.13) $(z_{1}-w+\hbar)a(z_{1})a(w)_{0}b(w)=(z_{1}-w+3\hbar)a(w)_{0}b(w)a(z_{1}).$
In view of Lemma 4.3, we have
$a(z_{1})a(w)_{0}b(w)-\!\left(\frac{w-z_{1}-3\hbar}{w-z_{1}-\hbar}\right)\!a(w)_{0}b(w)a(z_{1})=a(w)_{0}a(w)_{0}b(w)z_{1}^{-1}\delta\!\left(\frac{w-\hbar}{z_{1}}\right)\!,$
$a(z_{2})a(w)_{0}b(w)-\!\left(\frac{w-z_{2}-3\hbar}{w-z_{2}-\hbar}\right)\!a(w)_{0}b(w)a(z_{2})=a(w)_{0}a(w)_{0}b(w)z_{2}^{-1}\delta\!\left(\frac{w-\hbar}{z_{2}}\right)\!.$
Combining these relations with (4.12) (applying $\operatorname{Res}_{z_{1}}\operatorname{Res}_{z_{2}}$) we get
$a(w)_{0}a(w)_{0}b(w)=-a(w)_{0}a(w)_{0}b(w),$
which implies $a(w)_{0}a(w)_{0}b(w)=0$. Then we obtain
(4.14) $a(z)a(w)_{0}b(w)-\left(\frac{w-z-3\hbar}{w-z-\hbar}\right)a(w)_{0}b(w)a(z)=0.$
It is clear that this relation also implies (4.12). Now, the proof is complete. ∎
We continue to give another characterization of the Serre relation. First, we present the following two technical results:
@@
Let $a(x),b(x)\in{\mathcal{E}}_{\hbar}(W),\ \nu\in{\mathbb{C}}$ such that
(4.21) $(z-w-2\nu\hbar)a(z)a(w)=(z-w+2\nu\hbar)a(w)a(z),$
(4.22) $(z-w+\nu\hbar)a(z)b(w)=(z-w-\nu\hbar)b(w)a(z).$
Set
(4.23) ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}=(z-w+\nu\hbar)a(z)b(w),$
(4.24) ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{1})a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
$=\$$(z_{1}-z_{2})^{-1}(z_{1}-z_{2}-2\nu\hbar)(z_{1}-w+\nu\hbar)(z_{2}-w+\nu\hbar)a(z_{1})a(z_{2})b(w).$
@@
(4.25) ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z=w-\nu\hbar}=a(w)_{0}b(w),$

### v2

@@ Abstract.
For any symmetrizable generalized Cartan matrix $A$, we introduce an algebra $\widehat{\mathcal{DY}}(A)$, which is essentially the centrally extended double Yangian when $A$ is of finite type, and we give a new field (current) presentation of $\widehat{\mathcal{DY}}(A)$. Among the main results, for any $\ell\in{\mathbb{C}}$ we construct a universal vacuum $\widehat{\mathcal{DY}}(A)$-module $\mathcal{V}_{A}(\ell)$ of level $\ell$, prove that there exists a natural $\hbar$-adic weak quantum vertex algebra structure on $\mathcal{V}_{A}(\ell)$, and give an isomorphism between the category of restricted $\widehat{\mathcal{DY}}(A)$-modules of level $\ell$ and the category of $\mathcal{V}_{A}(\ell)$-modules.
@@ 1. Introduction
In a previous paper [JKLT2] joint with Jing and Tan, by using the smash product construction of nonlocal vertex algebras (see [Li4]), we constructed a family of $\hbar$-adic quantum vertex algebras by explicitly deforming lattice vertex algebras. It is remarkable that Drinfeld’s quantum affinization process can be extended to general symmetrizable quantum Kac-Moody algebras ([GKV, J, N, CJKT]). In [Kon], mainly by deforming affine vertex algebras the first named author of this paper constructed a family of $\hbar$-adic quantum affine vertex algebras and their $\phi$-coordinated quasi modules, based on Drinfeld presentations.
In this current paper, we continue studying centrally extended double Yangians in the context of $\hbar$-adic quantum vertex algebras and their modules, using the Drinfeld-like generator-and-defining-relation approach. We first introduce an algebra $\widehat{\mathcal{DY}}(A)$ for any symmetrizable generalized Cartan matrix $A$, which essentially coincides with the centrally extended double Yangian when $A$ is of finite type, and we then give a new field (current) presentation of $\widehat{\mathcal{DY}}(A)$. Among the main results, for any $\ell\in{\mathbb{C}}$, we construct a universal vacuum $\widehat{\mathcal{DY}}(A)$-module $\mathcal{V}_{A}(\ell)$ of level $\ell$, prove that there exists a natural $\hbar$-adic weak quantum vertex algebra structure on $\mathcal{V}_{A}(\ell)$, and give an isomorphism between the category of restricted $\widehat{\mathcal{DY}}(A)$-modules of level $\ell$ and the category of $\mathcal{V}_{A}(\ell)$-modules.
@@ 4. Vertex-operator Serre relation and iterate formula
Let $a(x),b(x)\in{\mathcal{E}}_{\hbar}(W)$, and $m\in{\mathbb{Z}}_{+}$. The relation
(4.7) $\sum_{\sigma\in S_{m}}\sum_{r=0}^{m}\binom{m}{r}(-1)^{r}a(z_{\sigma(1)})a(z_{\sigma(2)})\cdots a(z_{\sigma(r)})b(w)a(z_{\sigma(r+1)})\cdots a(z_{\sigma(m)})=0$
is referred to as the Serre-like relation for $(a(z),b(z),m)$.
We use a similar but simpler method to the proof of [CJKT, Theorem 5.16] to give another characterization of the Serre-like relation. First, we present the following two technical results:
@@
Let $a(x),b(x)\in{\mathcal{E}}_{\hbar}(W)$ and let $\mu\in{\mathbb{C}}^{\times}$, $\nu\in{\mathbb{C}}$ such that
(4.14) $(z-w-2\mu\hbar)a(z)a(w)=(z-w+2\mu\hbar)a(w)a(z),$
(4.15) $(z-w+\mu\nu\hbar)a(z)b(w)=(z-w-\mu\nu\hbar)b(w)a(z).$
For $r\in{\mathbb{N}}$, we set
${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{1})\cdots a(z_{r})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}=\prod_{1\leq i<j\leq r}\frac{z_{i}-z_{j}-2\mu\hbar}{z_{i}-z_{j}}\prod_{1\leq i\leq r}(z_{i}-w+\mu\nu\hbar)a(z_{1})\cdots a(z_{r})b(w).$
@@
(4.16) ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{1})\cdots a(z_{r})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}\in{\mathcal{E}}_{\hbar}^{(r+1)}(W):=\operatorname{Hom}(W,W_{\hbar}((z_{1},\dots,z_{r},w))),$
(4.17) $a(w)_{0}^{r}b(w)$
$\quad={\mathopen{\overset{\bullet}{{{\bullet}}}}{a(w+(2r-2-\nu)\mu\hbar)a(w+(2r-4-\nu)\mu\hbar)\cdots a(w-\nu\mu\hbar)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}},$
(4.18) $a(z)a(w)_{0}^{r}b(w)=\frac{w-z+\mu\nu\hbar}{w-z-\mu\nu\hbar}\prod_{s=1}^{r}\lim_{z_{s}\to w+(2r-2s-\mu\nu)\hbar}\frac{z_{s}-z-2\mu\hbar}{z_{s}-z+2\mu\hbar}a(w)_{0}^{r}b(w)a(z)$
$\quad+z^{-1}\delta\left(\frac{w-(\nu-2r)\mu\hbar}{z}\right)a(w)_{0}^{r+1}b(w).$
Proof.
It is no harm to assume $\mu=1$ Applying Lemma 4.5 to equation (4.14), we get
(4.19) $\frac{z-w-2\hbar}{z-w}a(z)a(w)=\frac{w-z-2\hbar}{w-z}a(w)a(z).$
Then the first relation follows immediate from (4.19) and (4.15). We next prove the relation (4.17) by using induction on $r$. If $r=0$, there is nothing need to prove. Suppose it holds for $r$. Then
$(z-w+\nu\hbar)\prod_{s=1}^{r}\lim_{z_{s}\to w+(2r-2s-\nu)\hbar}\frac{z-z_{s}-2\hbar}{z-z_{s}}a(z)a(w)_{0}^{r}b(w)$
$=$ ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)a(w+(2r-2-\nu)\hbar)a(w+(2r-4-\nu)\hbar)\cdots a(w-\nu\hbar)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}},$
$(-w+z-\nu\hbar)\prod_{s=1}^{r}\lim_{z_{s}\to w+(2r-2s-\nu)\hbar}\frac{z_{s}-z-2\hbar}{z_{s}-z}a(w)_{0}^{r}b(w)a(z)$
$=$ ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)a(w+(2r-2-\nu)\hbar)a(w+(2r-4-\nu)\hbar)\cdots a(w-\nu\hbar)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}.$
Notice that
$(z-w+\nu\hbar)\prod_{s=1}^{r}\lim_{z_{s}\to w+(2r-2s-\nu)\hbar}\frac{z-z_{s}-2\hbar}{z-z_{s}}=(z-w+(\nu-2r)\hbar).$
By Lemma 4.3, we get that
$z^{-1}\delta\left(\frac{w-(\nu-2r)\hbar}{z}\right)a(w)_{0}^{r+1}b(w)$
(4.20) $=$ $a(z)a(w)_{0}^{r}b(w)-\frac{w-z+\nu\hbar}{w-z-\nu\hbar}\prod_{s=1}^{r}\lim_{z_{s}\to w+(2r-2s-\nu)\hbar}\frac{z_{s}-z-2\hbar}{z_{s}-z+2\hbar}a(w)_{0}^{r}b(w)a(z)$
$=$ $z^{-1}\delta\left(\frac{w-(\nu-2r)\hbar}{z}\right)$
$\times{\mathopen{\overset{\bullet}{{{\bullet}}}}{a(w+(2r-\nu)\hbar)a(w+(2r-2-\nu)\hbar)\cdots a(w-\nu\hbar)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}.$
Taking $\operatorname{Res}_{z}$ on both hand sides, we get that
$a(w)_{0}^{r+1}b(w)={\mathopen{\overset{\bullet}{{{\bullet}}}}{a(w+(2r-\nu)\hbar)a(w+(2r-2-\nu)\hbar)\cdots a(w-\nu\hbar)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}},$
as desired. The last relation follows immediate from the relation (4.20). ∎
We introduce some notations and conventions. For a set $A$, we denote by $|A|$ the cardinal of $A$. For two integers $m\leq n$, we denote by $[m,n]$ the set $\left.\left\{{k\in{\mathbb{Z}}}\,\right|\,{m\leq k\leq n}\right\}$}. Given a natural number $r\leq n-m$ and a subset $I\subset[m,n]$ of cardinal $r$, we write
$I={i_{1}<i_{2}<\cdots<i_{r}},\quad[0,n]\setminus I={i^{\prime}_{1}<i^{\prime}_{2}<\cdots<i^{\prime}_{n-r}}.$
Lemma 4.7.
Let $a(x),b(x)\in{\mathcal{E}}_{\hbar}(W)$ and let $\mu\in{\mathbb{C}}^{\times}$, $\nu\in{\mathbb{C}}$ such that (4.14) and (4.15) hold. Then for any $r\in{\mathbb{N}}$, we have that
$a(z_{1})a(z_{2})\cdots a(z_{r})b(w)=\sum_{I\subset[1,r]}\prod_{s=1}^{r-|I|}\bigg(\frac{w-z_{i^{\prime}_{s}}+\mu\nu\hbar}{w-z_{i^{\prime}_{s}}-\mu\nu\hbar}\prod_{\begin{subarray}{c}1\leq t\leq|I|\\ i^{\prime}_{s}<i_{t}\end{subarray}}\frac{z_{i_{t}}-z_{i^{\prime}_{s}}-2\mu\hbar}{z_{i_{t}}-z_{i^{\prime}_{s}}+2\mu\hbar}\bigg)$
$\times a(w)_{0}^{|I|}b(w)a(z_{i^{\prime}_{1}})a(z_{i^{\prime}_{2}})\cdots a(z_{i^{\prime}_{r-|I|}})\prod_{t=1}^{|I|}z_{i_{t}}^{-1}\delta\left(\frac{w-(\nu-2(|I|-t))\mu\hbar}{z_{i_{t}}}\right).$
Proof.
It is no harm to assume $\mu=1$. If $r=0$, then there is nothing need to prove. We suppose this lemma holds true for $r$. From Lemma 4.6
$a(z_{0})a(z_{1})a(z_{2})\cdots a(z_{r})b(w)$
$=$ $\sum_{I\subset[1,r]}\prod_{s=1}^{r-|I|}\bigg(\frac{w-z_{i^{\prime}_{s}}+\nu\hbar}{w-z_{i^{\prime}_{s}}-\nu\hbar}\prod_{\begin{subarray}{c}1\leq t\leq|I|\\ i^{\prime}_{s}<i_{t}\end{subarray}}\frac{z_{i_{t}}-z_{i^{\prime}_{s}}-2\hbar}{z_{i_{t}}-z_{i^{\prime}_{s}}+2\hbar}\bigg)$
$\times a(z_{0})a(w)_{0}^{|I|}b(w)a(z_{i^{\prime}_{1}})a(z_{i^{\prime}_{2}})\cdots a(z_{i^{\prime}_{r-|I|}})\prod_{t=1}^{|I|}z_{i_{t}}^{-1}\delta\left(\frac{w-(\nu-2(|I|-t))\hbar}{z_{i_{t}}}\right)$
$=$ $\sum_{I\subset[1,r]}\prod_{s=1}^{r-|I|}\bigg(\frac{w-z_{i^{\prime}_{s}}+\nu\hbar}{w-z_{i^{\prime}_{s}}-\nu\hbar}\prod_{\begin{subarray}{c}1\leq t\leq|I|\\ i^{\prime}_{s}<i_{t}\end{subarray}}\frac{z_{i_{t}}-z_{i^{\prime}_{s}}-2\hbar}{z_{i_{t}}-z_{i^{\prime}_{s}}+2\hbar}\bigg)$
$\times a(w)_{0}^{|I|+1}b(w)a(z_{i^{\prime}_{1}})a(z_{i^{\prime}_{2}})\cdots a(z_{i^{\prime}_{r-|I|}})\prod_{t=1}^{|I|}z_{i_{t}}^{-1}\delta\left(\frac{w-(\nu-2(|I|-t))\hbar}{z_{i_{t}}}\right)$
$\times z_{0}^{-1}\delta\left(\frac{w-2|I|\hbar}{z_{0}}\right)$
$+$ $\sum_{I\subset[1,r]}\prod_{s=1}^{r-|I|}\bigg(\frac{w-z_{i^{\prime}_{s}}+\nu\hbar}{w-z_{i^{\prime}_{s}}-\nu\hbar}\prod_{\begin{subarray}{c}1\leq t\leq|I|\\ i^{\prime}_{s}<i_{t}\end{subarray}}\frac{z_{i_{t}}-z_{i^{\prime}_{s}}-2\hbar}{z_{i_{t}}-z_{i^{\prime}_{s}}+2\hbar}\bigg)\times\frac{w-z_{0}+\nu\hbar}{w-z_{0}-\nu\hbar}\prod_{t=1}^{|I|}\frac{z_{i_{t}}-z_{0}-2\hbar}{z_{i_{t}}-z_{0}+2\hbar}$
$\times a(w)_{0}^{|I|}b(w)a(z_{0})a(z_{i^{\prime}_{1}})a(z_{i^{\prime}_{2}})\cdots a(z_{i^{\prime}_{r-|I|}})\prod_{t=1}^{|I|}z_{i_{t}}^{-1}\delta\left(\frac{w-(\nu-2(|I|-t))\hbar}{z_{i_{t}}}\right)$
$=$ $\sum_{I\subset[0,r]}\prod_{s=1}^{r+1-|I|}\bigg(\frac{w-z_{i^{\prime}_{s}}+\nu\hbar}{w-z_{i^{\prime}_{s}}-\nu\hbar}\prod_{\begin{subarray}{c}1\leq t\leq|I|\\ i^{\prime}_{s}<i_{t}\end{subarray}}\frac{z_{i_{t}}-z_{i^{\prime}_{s}}-2\hbar}{z_{i_{t}}-z_{i^{\prime}_{s}}+2\hbar}\bigg)$
$\times a(w)_{0}^{|I|}b(w)a(z_{i^{\prime}_{1}})a(z_{i^{\prime}_{2}})\cdots a(z_{i^{\prime}_{r-|I|}})\prod_{t=1}^{|I|}z_{i_{t}}^{-1}\delta\left(\frac{w-(\nu-2(|I|-t))\hbar}{z_{i_{t}}}\right),$
as desired. ∎
For a positive integer $m$, we denote by $S_{m}$ the symmetric group on $[1,m]$. Given an integer $p\in[0,m]$, we view $S_{p}$ as a subgroup of $S_{m}$ in the obvious way and introduce the following subgroups of $S_{m}$:
$S_{m}^{p}:={\left.\left\{{\sigma\in S_{m}}\,\right|\,{\sigma([p+1,m])\subset[p+1,m],\,\,\textnormal{{and}}\,\,\sigma(r)=r,\,\,\textnormal{{for }}r\leq p}\right\}},$
$S_{m,p}:={\left.\left\{{\sigma\in S_{m}}\,\right|\,{\sigma(s)<\sigma(t),\,\,\textnormal{{for }}s<t\leq p\,\,\textnormal{{or }}p<s<t}\right\}}.$
Then we have that
Lemma 4.8.
Let $a(x),b(x)\in{\mathcal{E}}_{\hbar}(W)$ and let $\mu\in{\mathbb{C}}^{\times}$, $\nu\in{\mathbb{N}}$ such that (4.14) and (4.15) hold. Set $m=\nu+1$. Then
$\sum_{\tau\in S_{m}}\sum_{r=0}^{m}\binom{m}{r}(-1)^{r}a(z_{\tau(1)})\cdots a(z_{\tau(r)})b(w)a(z_{\tau(r+1)})\cdots a(z_{\tau(m)})$
$=$ $\sum_{\tau\in S_{m}}\sum_{p=0}^{m}a(w)_{0}^{p}b(w)a(z_{\tau(p+1)})a(z_{\tau(p+2)})\cdots a(z_{\tau(m)})\prod_{s=1}^{p}z_{\tau(s)}^{-1}\delta\left(\frac{w-(\nu-2(p-s))\mu\hbar}{z_{\tau(s)}}\right)$
$\times\sum_{\sigma\in S_{m,p}}\sum_{r=\sigma(p)}^{m}\binom{m}{r}(-1)^{r}\prod_{s=p+1}^{r}\frac{w-z_{\tau(s)}+\mu\nu\hbar}{w-z_{\tau(s)}-\mu\nu\hbar}\prod_{\begin{subarray}{c}1\leq s\leq p<t\leq r\\ \sigma(s)>\sigma(t)\end{subarray}}\frac{z_{\tau(s)}-z_{\tau(t)}-2\mu\hbar}{z_{\tau(s)}-z_{\tau(t)}+2\mu\hbar}.$
Proof.
It is no harm to assume that $\mu=1$. For a subset $I\subset[1,m]$, we set $m(I)=\max I$ if $I\neq\emptyset$ or 0 if $I=\emptyset$. From Lemma 4.7, we have that
$\sum_{r=0}^{m}\binom{m}{r}(-1)^{r}a(z_{1})\cdots a(z_{r})b(w)a(z_{r+1})\cdots a(z_{m})$
$=$ $\sum_{r=0}^{m}\binom{m}{r}(-1)^{r}\sum_{I\subset[1,r]}\prod_{s=1}^{r-|I|}\bigg(\frac{w-z_{i^{\prime}_{s}}+\nu\hbar}{w-z_{i^{\prime}_{s}}-\nu\hbar}\prod_{\begin{subarray}{c}1\leq t\leq|I|\\ i^{\prime}_{s}<i_{t}\end{subarray}}\frac{z_{i_{t}}-z_{i^{\prime}_{s}}-2\hbar}{z_{i_{t}}-z_{i^{\prime}_{s}}+2\hbar}\bigg)$
$\times a(w)_{0}^{|I|}b(w)a(z_{i^{\prime}_{1}})a(z_{i^{\prime}_{2}})\cdots a(z_{i^{\prime}_{r-|I|}})a(z_{r+1})\cdots a(z_{m})$
$\times\prod_{s=1}^{|I|}z_{i_{s}}^{-1}\delta\left(\frac{w-(\nu-2(|I|-s))\hbar}{z_{i_{s}}}\right)$
$=$ $\sum_{I\subset[1,m]}\sum_{r=m(I)}^{m}\binom{m}{r}(-1)^{r}a(w)_{0}^{|I|}b(w)a(z_{i^{\prime}_{1}})a(z_{i^{\prime}_{2}})\cdots a(z_{i^{\prime}_{m-|I|}})$
$\times\prod_{s=1}^{r-|I|}\bigg(\frac{w-z_{i^{\prime}_{s}}+\nu\hbar}{w-z_{i^{\prime}_{s}}-\nu\hbar}\prod_{\begin{subarray}{c}1\leq t\leq|I|\\ i^{\prime}_{s}<i_{t}\end{subarray}}\frac{z_{i_{t}}-z_{i^{\prime}_{s}}-2\hbar}{z_{i_{t}}-z_{i^{\prime}_{s}}+2\hbar}\bigg)\prod_{s=1}^{|I|}z_{i_{s}}^{-1}\delta\left(\frac{w-(\nu-2(|I|-s))\hbar}{z_{i_{s}}}\right)$
$=$ $\sum_{p=0}^{m}\sum_{\sigma\in S_{m,p}}\sum_{r=\sigma(p)}^{m}\binom{m}{r}(-1)^{r}a(w)_{0}^{p}b(w)a(z_{\sigma(p+1)})a(z_{\sigma(p+2)})\cdots a(z_{\sigma(m)})$
$\times\prod_{s=p+1}^{r}\frac{w-z_{\sigma(s)}+\nu\hbar}{w-z_{\sigma(s)}-\nu\hbar}\prod_{\begin{subarray}{c}1\leq s\leq p<t\leq r\\ \sigma(s)>\sigma(t)\end{subarray}}\frac{z_{\sigma(s)}-z_{\sigma(t)}-2\hbar}{z_{\sigma(s)}-z_{\sigma(t)}+2\hbar}$